OpenStudy (anonymous):

4x – (6 – x) = 12 + 4x

OpenStudy (midhun.madhu1987):

4x – (6 – x) = 12 + 4x 4x - 6 + x = 12 + 4x 4x + x - 4x = 12 + 6 Find x.